Show that the space R[a,b] of all Riemann integrable functions on the interval [a,b] is a linear space over a vector field R


1
Expert's answer
2021-11-01T13:18:37-0400

"\int((f+g)(x))dx=\int f(x)dx+\int g(x)dx"


"\int(af(x))dx=a\int f(x)dx"


So, space R[a,b] of all Riemann integrable functions is a linear space.